OpenStudy (anonymous):

Please help me! Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (-8, 0), (1, 5)

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

Slope of line through the two points is: \[m=\frac{ y _{2} -y _{1} }{ x _{2} -x _{1} }=\frac{ 5-0 }{ 1-(-8) }=\frac{ 5 }{ 9 }\] Using the point slope form of the equation of a line: \[y-y_{1}=m(x-x_{1}) \rightarrow y-0=\frac{ 5 }{ 9 }(x+8 )\] Multiply both sides of the equation by 9 and rearranging: \[9y=5x+40 \rightarrow -5x+9y=40\]